OpenStudy (anonymous):

An angle bisector of a triangle divides the opposite side of the triangle into segments 6 cm and 5 cm long. A second side of the triangle is 6.9 cm long. Find the longest and shortest possible lengths of the third side of the triangle. Round answers to the nearest tenth of a centimeter.

OpenStudy (ash2326):

Do you have a figure?

OpenStudy (anonymous):

there's no figure. /: sorry.

OpenStudy (ash2326):

No worries, that's why it has asked for two answers. Let ABC be the triangle and AD be the angle bisector. Let BD=6 and DC=5  Do you understand the figure?

OpenStudy (anonymous):

yes!

OpenStudy (ash2326):

@abm1995 ??

OpenStudy (anonymous):

do you want the choices?

OpenStudy (anonymous):

a. 41.4 cm, 8.3 cm b. 30 cm, 5.8 cm c. 41.4 cm, 4.3 cm d. 8.3 cm, 5.8 cm

OpenStudy (anonymous):

does anyone understand this?

OpenStudy (ash2326):

Now we can have either AB or AC as 6.9 cm, we don't know that. It's the second side. Do you know Angle Bisector theorem?

OpenStudy (anonymous):

no i don't./: i'm dumb.

OpenStudy (anonymous):

do you know the answer? then when i figure it out you will say if i'm right or not?

OpenStudy (ash2326):

Yeah, I'll do let me explain you angle bisector theorem

OpenStudy (anonymous):

ok.

OpenStudy (ash2326):

Angle bisector divides the side opposite to the angle in the same ratio as that of the other two sides of the triangle, for example here \[\frac{BD}{DC}=\frac {AB}{AC}\]

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

wouldn't the answer be B then?

OpenStudy (ash2326):

here we have BD=6 DC=5 Now assume AB as 6.9, find AC ???

OpenStudy (anonymous):

so b?

OpenStudy (ash2326):

No not yet. We have to find one more thing, assume AC=6.9 and find AB

OpenStudy (anonymous):

oh it'd be c then or a.

OpenStudy (ash2326):

What did you get for AB?

OpenStudy (anonymous):

8.3?

OpenStudy (ash2326):

Yes, we don't know which is the second side AB or AC. So we can get either 8.3 or 5.8

OpenStudy (anonymous):

so , d?

OpenStudy (ash2326):

yes, did you understand?

OpenStudy (anonymous):

yes! your a lot of help!

OpenStudy (ash2326):

No Problem :)

OpenStudy (anonymous):

i need help with a little more../: i'm behind in school. could you help me?

OpenStudy (anonymous):

What is the value of x?

OpenStudy (anonymous):

a. 5 b. 2.5 c. 7.5 d. 10

OpenStudy (ash2326):

The two lines indicated with arrow are parallel, so the two triangles are similar

OpenStudy (anonymous):

yes. i know.

OpenStudy (ash2326):

\[\frac{x}{x+x+5}=\frac {x-2}{x-2+x+1}\] Now solve for x

OpenStudy (anonymous):

wouldn't it be 5 though?

OpenStudy (ash2326):

yes, just plugin the value and check

OpenStudy (anonymous):

ok, so my answer would be a 5?

OpenStudy (ash2326):

check it, you should ascertain it yourself

OpenStudy (anonymous):

i don't think it's 5.

OpenStudy (ash2326):

What did you get when you plugged in x=5 ?

OpenStudy (anonymous):

idk. i'm just trying to get this done. i'm so behind.

OpenStudy (ash2326):

Relax and do this \[\frac{x}{x+x+5}=\frac {x-2}{x-2+x+1}\] \[\frac {x}{2x+5}=\frac{x-2}{2x-1}\] Put x=5 and check

OpenStudy (anonymous):

i got 1/3

OpenStudy (ash2326):

Both sides?

OpenStudy (anonymous):

yes. do you know the answer?

OpenStudy (ash2326):

So 5 is the correct answer

OpenStudy (anonymous):

do you know that it is?

OpenStudy (ash2326):

yes, that's why we got the same thing on both sides. Check other options also, you'll see that the left side won't be equal to the right side

OpenStudy (anonymous):

What similarity statement can you write relating the three triangles in the diagram below?

OpenStudy (anonymous):